# Jinhao Zhang1  Yunquan Zhang2  Boyang Zhang3,4  Zeyu Liu5  Daning Cheng2

### v1

@@ Introduction
Quantization techniques exhibit notable advantages in compressing deep neural network models and minimizing computational overhead. However, their practical deployment faces several critical challenges. Quantization inherently involves mapping high-precision floating-point parameters to low-bit representations, a process that inevitably introduces information loss, thereby degrading model generalization and accuracy. This degradation is particularly pronounced under ultra-low bit-width settings (e.g., below 8 bits), where quantization noise can significantly impair representational capacity, leading to performance deterioration—especially in complex tasks or applications with stringent accuracy requirements. Thus, achieving high efficiency without compromising model performance remains a central research challenge.
The MoE (Mixture of Experts) system is an artificial intelligence architecture based on a combination of expert models, designed to solve complex problems by integrating the capabilities of multiple specialized models (the ”experts”). Each expert handles a specific aspect of the input data, while a component called the ”gating network” determines how to allocate tasks to different experts based on the input’s characteristics. We named gating network as router model in this paper. This approach allows the model to scale to very large sizes while maintaining computational efficiency, as not all experts need to compute every input. MoE systems are particularly suitable for applications requiring high flexibility and strong expressive power, such as NLP and image recognition, delivering more accurate and personalized services and predictions. As technology advances, MoE has become a key technique for building efficient, large-scale machine learning models. However, there are few works which focus on how to improve quantization model performance via MoE system.
To mitigate the performance degradation caused by model quantization, make full use of computing resources, and avoid significantly decreasing the inference speed of the quantization model, we propose Mixture of Quantization Expert methods (MoQE)—an expert system for quantization inference based on the MoE paradigm. Prior studies (Liu et al. 2025) indicate that different quantization variants of the same full-precision model exhibit heterogeneous performance degradation across distinct sub-datasets. Motivated by this observation, MoQE assigns different sub-datasets to specialized quantization models—termed ”quantization experts”—thereby improving the overall system performance. Notably, as the number of quantization experts increases, the coverage and representational diversity of the system expand, leading to progressive performance gains.
@@
Our contributions are listed as the following:
1.We design a Mixture of Quantized Experts system, abbr.MoQE. By integrating quantization models into an expert system, MoQE routes different data subsets to the quantization models where they perform best, resulting in overall system performance superior to that of any single expert.
2.We design data router models for both CV and NLP tasks. The router model is a lightweight classification model with a short inference time, achieving inference speed of MoQE system comparable to that of a single quantization model.
@@
3.Through experiments, we have validated the performance of MoQE on both CV and NLP tasks. Experimental results show that the performance of the MoQE system exceeds that of any individual quantization expert model comprising it. Furthermore, the performance of the MoQE system improves as the number of expert models increases. And experiments code address is shown in the Appendix.
Related Works
Quantization is one of the most popular techniques for compressing neural networks. It can be roughly divided into two main methods: quantization-aware training (QAT) and post-training quantization (PTQ). QAT considers quantization during the network training phase, while PTQ quantizes the network after training. PTQ is widely used in network deployment due to its low time consumption and computational resource requirements. HAWQ-v2 (Dong et al. 2019) uses the trace of the layer weight’s Hessian matrix as the index to show this layer’s sensitivity. The work (Zhe et al. 2019) formulates the mixed-precision problem of weights and activations as a Lagrangian optimization problem whose solution is the optimum joint precision allocation for all weights and activations. AdaQuant (Hubara et al. 2020) optimizes its parameters on the calibration set to minimize the quantization error of each layer separately. (Liu et al. 2023) alleviate the overfitting problem in PTQ caused by the small number of calibration sets by adjusting the distribution of activations. (Shang et al. 2024) introduced mutual information into PTQ calibration to optimize the quantization parameters. These methods focus on the quantization parameters to alleviate the quantization error, but cannot achieve good results at very low bits and run slowly. Our method uses performance and speed improvements at very low bits, and does not require calibration sets and fine-tuning.
Mixture of Experts (MoE) (Jacobs et al. 1991; Eigen, Ranzato, and Sutskever 2013) is a composite modeling framework comprising multiple specialized sub-models, referred to as experts, which are combined through a structured integration mechanism. There are two kinds of MoE paradigm hard routers and soft router. In the hard routing paradigm, experts are typically pre-assigned to handle specific, predefined patterns or modalities. This design choice is motivated by the inherent disparities among multi-modal data (Liang et al. 2022), which can hinder soft routing mechanisms from effectively learning optimal token-to-expert assignment policies. A number of studies (Bao et al. 2022; Long et al. 2023; Shen et al. 2023) exploit this modality-level separation by explicitly decoupling experts according to data modalities, with each expert dedicated to processing a single modality. A key advantage of such hard routing approaches is that they eliminate the need for training or learning the router, simplifying deployment. This strategy has also been widely adopted in task-specific MoE architectures (Li et al. 2023; Ma et al. 2023). In contrast, soft routers have been extensively explored in NLP (Shazeer et al. 2017; Fedus, Zoph, and Shazeer 2022; Komatsuzaki et al. 2022; Zoph et al. 2022), where routing decisions are made probabilistically, enabling dynamic and context-aware allocation of tokens to different experts. This flexibility allows each expert to specialize in particular semantic or syntactic patterns, promoting model sparsity and scalability. Our work primarily focuses on leveraging soft routing mechanisms within the MoE framework. Notably, soft-routed MoE architectures have also been adapted to multimodal learning in smaller-scale models (e.g., million-parameter range), such as EVE (Chen et al. 2024) and LIMoE (Mustafa et al. 2022), which utilize soft routers to facilitate cross-modal fusion.
Methodological
Mixture of Quantization Experts
@@
To mitigate the degradation of model accuracy caused by quantization, this paper proposes a novel framework termed Mixture of Quantization Experts (MoQE), the working pattern of MoQE is shown in Figure  1. In the MoQE architecture, multiple quantization variants of the same full-precision model are integrated through an expert system paradigm. These quantization models are typically derived via different quantization methods on the same full-precision model. The selection of which specific quantization model to activate is governed by a dedicated router model. The design and formulation of the router model will be detailed in the following section. The activated quantization expert model performs the inference on this sample. The router model is trained by a train dataset.
@@
Figure 1: The working pattern of MoQE. The data sample first passes through a router model which may involve embedding process, to determine which quantization expert model to activate, and then is fed into the corresponding quantization expert model for inference.
@@
The theoretical foundation of this approach stems from the observation that different quantization models exhibit distinct biases (unfairness) across various regions of the input data space (Liu et al. 2025). Specifically, for certain subsets of data, one quantization model may consistently yield lower prediction errors (i.e., lower loss values) than others. By leveraging the router model to dynamically assign samples to the most suitable expert, MoQE effectively selects, at each inference step, the model with the lowest loss function value on this subset of data. As illustrated in the Figure  2, this strategy enables the system to synthesize a composite decision boundary that outperforms any individual quantization model. Consequently, the router model can be interpreted as a multi-class classifier, and its effective design is critical to ensuring that the overall system performance exceeds that of any constituent model.
@@
Figure 2: The explanations of MoQE system which consists of two quantization experts. MoQE always selects the model with the lower loss on this subset of data.
@@
It is important to note that the router model operates on vectorized representations of input data, implying that raw data formats may first be transformed into appropriate embedding. This process is optional depending on the tasks. For example, in CV tasks, images are naturally represented as tensors and can be directly processed by the router model. Thus, CV tasks do not need embedding process. In contrast, for NLP tasks, textual inputs must be converted into dense vector representations—typically through embedding techniques—before they can be handled. The quality of this embedding process directly influences the router model’s classification accuracy.
@@
In the context of NLP, MoQE uses the pre-existing embedding layer of the original full-precision model as its embedding layer. This choice serves two purposes: (1) it avoids the need to train a separate embedding layer for routing, which would be computationally expensive and potentially degrade performance due to misalignment; and (2) reducing inference cost, using the same embedding layer saves one inference step through the embedding layer. Therefore, the embedding layer of the quantization expert models is not quantized.
@@
In our experimental design, we primarily focus on quantization variants of a single full-precision model, reflecting typical industrial deployment scenarios where only one full-precision model is available for quantization. Moreover, in NLP tasks, the shared embedding layer constraint limits the framework to experts originating from the same base model, which limits that all quantization expert models are from the same full-precision model. However, the MoQE framework is, in principle, extensible to quantization models derived from multiple distinct full-precision models. Prior work (Liu et al. 2025) suggests that for well-trained models, performance variations across quantization versions in different data regions are largely determined by local gradient and Hessian matrix properties.
Router Model
@@
The design of router models must be tailored to the intrinsic characteristics of different task domains. For instance, NLP tasks necessitate a strong emphasis on contextual dependencies, while CV tasks primarily rely on spatial and hierarchical visual features. By carefully calibrating the architectural and parametric configurations of the router model, classification accuracy can be significantly improved, ensuring that each quantization expert model operates at peak efficacy within its targeted data distribution. This, in turn, contributes to enhanced overall inference accuracy.
Accordingly, in the development of task-specific routing architectures, we advocate for the reconstruction of foundational components based on domain-informed models. This strategy ensures heightened sensitivity and adaptability to the unique statistical and semantic properties of the input data within a given modality. To this end, and to address the prevalent use cases in modern deep learning, we have designed two specialized router models tailored specifically for CV and NLP tasks, respectively—each optimized to exploit the structural priors inherent in its respective domain.
@@
We present a lightweight, multi-head ResNet-8 router model for computer vision (CV) tasks that resolves the load-imbalance and computational redundancy endemic to conventional multi-expert architectures. The pipeline begins with an MLP that compresses a 2048-D global feature into $1024\to 128\to 1024$ dimensions, each followed by batch normalization (BN) and ReLU activation. The resulting 1024-D vector is reshaped into a $48\times 16\times 8\times 8$ tensor, preserving fine-grained spatial cues while remaining amenable to efficient convolution. A three-stage SEResNet-8 (with channel dimensions 16/32/64, stride pattern $1\to 2\to 2$, and one BasicBlock per stage) then extracts local representations. Finally, the feature map is flattened into a sequence and fed to an 8-head self-attention module (head dimension 8, learnable temperature $\tau=1.0$) to yield the routing logits, ensuring globally-aware yet balanced expert assignment with minimal overhead.
NLP Router Model
The core design principle of our router model in the NLP Router is to endow routing decisions with the same level of semantic richness and contextual sensitivity as the expert computations themselves. Conventional router model typically apply only a single linear transformation to token embeddings, which inherently restricts their receptive field to a local, context-agnostic perspective. This limitation often results in ambiguous or suboptimal routing, as tokens—deprived of their broader linguistic context—may be misclassified. To overcome this, we propose a deep, hierarchical context-aware router that systematically enriches token representations through multiple stages of contextual processing. The structure of router model is shown in Figure  3.
Figure 3: The structure of router model
[…]
Subsequent to this initial contextual encoding, we introduce a dedicated self-attention module designed for context refinement and selective focus. This module comprises independent linear projections for query, key, and value (each of dimension embed_dim), with attention scores scaled by $\sqrt{\mathrm{embed\_dim}/\mathrm{num\_heads}}$. It further incorporates attention dropout (attn_drop=0.1), a projection layer (proj), projection dropout (proj_drop=0.1), and a final layer normalization (norm). Far from being redundant, this stage functions as a context-aware refinement mechanism, enabling the model to dynamically re-weight the relevance of different contextual elements in determining the optimal routing path for each token.
The resulting representations are then passed through a final multilayer perceptron (MLP) to compute routing logits. This MLP consists of a dimensionality reduction layer ($\mathrm{embed\_dim}\to\mathrm{embed\_dim}/2$), followed by GELU activation, dropout, and a linear projection to the number of experts ($\mathrm{num\_experts}$). This multi-stage processing pipeline—encompassing global contextual encoding, context refinement via self-attention, and non-linear score generation—ensures that routing decisions are grounded in a comprehensive and fine-grained understanding of the input’s semantic intent.
Nevertheless, a complex router model necessitates a robust stabilization framework to prevent pathological behaviors such as expert collapse, a common failure mode in which only a small subset of experts is consistently activated. To mitigate this, we implement a comprehensive, multi-faceted load-balancing strategy comprising three key components:
  1. 1.
    An explicit load-balancing loss term that encourages uniform expert utilization within each batch;
  2. 2.
    Dynamic, history-aware bias adjustment, wherein expert usage frequencies are tracked via an exponential moving average (EMA, decay factor $\alpha=0.8$), and a negative bias proportional to utilization is applied to discourage overuse (scaled by a balance strength coefficient of 0.8);
  3. 3.
    Controlled noise injection during training, involving soft additive noise governed by a learnable parameter (noise_epsilon=1.0), combined with temperature-scaled softmax ($\mathrm{gating\_temperature}=1.0$). This noise is gradually annealed during training to balance exploration and convergence.
MoQE and Parallel Inference
In the broader context of inference frameworks, particularly beyond the scope of quantization, a critical challenge emerges: simply scaling up the number of GPUs does not always lead to a faster inference speed when model is inferring a single sample. Unlike traditional parallel computing paradigms, where increased hardware resources typically translate to faster execution, inference workloads often fail to benefit proportionally from additional GPUs. This is primarily because the inference process is constrained by the model’s architectural design and its inherent degree of parallelism. As a result, adding more GPUs can lead to underutilization of computational resources. Furthermore, the overhead associated with inter-device communication and synchronization can exacerbate inefficiencies, potentially degrading overall system performance rather than improving it.
In contrast to the conventional paradigm of trading computational resources for reduced inference latency, MoQE introduces a novel approach that leverages increased computational resources to improve inference quality—manifested as lower PPL in NLP tasks or higher classification accuracy in CV tasks. By adopting the MoE deployment strategy employed in DeepSeek, distinct quantization experts can be distributed across different computational devices. DeepSeek MoE deployment strategy keeps low inference latency. As our analysis demonstrates, the integration of a growing number of quantization expert models for specific data subsets enhances overall system performance while simultaneously enabling more effective utilization of hardware resources. This scalability in model experts offers a promising direction for achieving higher accuracy without compromising inference speed.
In addressing large-scale parallel inference tasks, conventional approaches rely on data parallelism by replicating a single quantization model across devices. This strategy effectively scales throughput through hardware augmentation.
The MoQE system presents a paradigm shift by introducing a quantization expert ensemble architecture with dynamic routing. Three key innovations enable better performance than data parallelism:
Decoupled Execution: Independent operation of quantization experts eliminates model-sharing bottlenecks, achieving data-parallel-like throughput through true parallelism.
Lightweight Routing: Sample-specific routing decisions incur minimal overhead while maintaining independence across processing units. When the same router model infers different samples, the inference processes are independent. Therefore, data parallelism can be applied to the routing model’s inference without affecting overall performance.
Experts-Driven Optimization: By assigning each sample to the most suitable expert based on data characteristics, the system maximizes model performance without compromising inference speed.
A workload-balanced MoQE system not only matches the throughput of data parallelism but also outperforms it in model accuracy. This dual advantage stems from its ability to simultaneously optimize for both computational efficiency and system performance.
Experiments

### v2

@@ 1 Introduction
Quantization techniques exhibit notable advantages in compressing deep neural network models and minimizing computational overhead. However, their practical deployment faces several critical challenges. Quantization inherently involves mapping high-precision floating-point parameters to low-bit representations, a process that inevitably introduces information loss, thereby degrading model generalization and accuracy. This degradation is particularly pronounced under ultralow bit-width settings (e.g., below 8 bits), where quantization noise can significantly impair representational capacity, leading to performance deterioration, especially in complex tasks or applications with stringent accuracy requirements. Thus, achieving high efficiency without compromising model performance remains a central research challenge.
The MoE (Mixture of Experts) system is an artificial intelligence architecture based on a combination of expert models, designed to solve complex problems by integrating the capabilities of multiple specialized models (the ”experts”). Each expert handles a specific aspect of the input data, while a component called the ”gating network” determines how to allocate tasks to different experts based on the input’s characteristics. In this paper, we named the gating network as a router model. This approach allows the model to scale to very large sizes while maintaining computational efficiency, as not all experts need to compute every input. MoE systems are particularly suitable for applications requiring high flexibility and strong expressive power, such as NLP and image recognition, delivering more accurate and personalized services and predictions. As technology advances, MoE has become a key technique for building efficient, large-scale machine learning models. However, there are few works which focus on how to improve quantization model performance via MoE system.
To address the common industrial constraint that only a single pretrained model is available for quantization, we propose the Mixture of Quantization Experts (MoQE), an MoE-style inference system that improves accuracy while keeping inference latency comparable to a single quantization model. Prior studies Liu et al. (2025) indicate that different quantization variants of the same full-precision model exhibit heterogeneous performance degradation across distinct sub-datasets. Motivated by this observation, MoQE assigns different sub-datasets to specialized quantization models—termed ”quantization experts”—thereby improving the overall system performance. Notably, as the number of quantization experts increases, the coverage and representational diversity of the system expand, leading to progressive performance gains.
@@
Our contributions are listed as the following: 1.We design a Mixture of Quantization Experts system, abbr. MoQE. By integrating quantization models into an expert system, MoQE routes different data subsets to the quantization models where they perform best, resulting in overall system performance superior to that of any single expert. 2.We design data router models for both CV and NLP tasks. The router model is a lightweight classification model with a short inference time, achieving inference speed of MoQE system comparable to that of a single quantization model. 3.Through experiments, we have validated the performance of MoQE on both CV and NLP tasks. Experimental results show that the performance of the MoQE system exceeds that of any individual quantization expert model comprising it. Furthermore, the performance of the MoQE system improves as the number of expert models increases. Experiments code address is https://github.com/paper-submission-goii/MoQE.
2 Related Work
Quantization falls into quantization-aware training (QAT) and post-training quantization (PTQ), with PTQ favored for deployment due to low cost. HAWQ-v2 Dong et al. (2019) uses Hessian trace to measure layer sensitivity, while Zhe et al. (2019) formulates mixed-precision allocation as Lagrangian optimization. AdaQuant Hubara et al. (2020) minimizes per-layer quantization error on calibration data. Recent works Liu et al. (2023); Shang et al. (2024) improve PTQ via activation distribution adjustment and mutual information, but still struggle at very low bits and require calibration. Our method avoids fine-tuning and calibration, achieving strong performance and efficiency in ultralow-bit regimes. Mixture of Experts(MoE) Jacobs et al. (1991); Eigen et al. (2013) employs multiple experts with routing mechanisms. Hard routing assigns experts to predefined modalities Bao et al. (2022); Long et al. (2023), simplifying deployment without learned routers. In contrast, soft routing in NLP Shazeer et al. (2017); Fedus et al. (2022) enables dynamic, probabilistic token-expert assignment for sparsity and scalability. Soft-routed MoEs are also adapted to multimodal models like EVE Chen et al. (2024) and LIMoE Mustafa et al. (2022). Our work focuses on soft routing for flexible, data-driven expert selection.
3 Methodology
3.1 Mixture of Quantization Experts
@@
To mitigate the degradation of model accuracy caused by quantization, this paper proposes a novel framework termed Mixture of Quantization Experts (MoQE), the working pattern of MoQE is shown in Figure  1. In the MoQE architecture, multiple quantization variants of the same full-precision model are integrated through an expert system paradigm. These quantization models are typically derived via different quantization methods on the same full-precision model. The selection of which specific quantization model to activate is governed by a dedicated router model. The design and formulation of the router model will be detailed in the following section. The activated quantization expert model performs the inference on this sample. The router is trained on a training dataset.
@@
Figure 1: The working pattern of MoQE. The data sample first passes through a router model which may involve an embedding step, to determine which quantization expert model to activate, and then is fed into the corresponding quantization expert model for inference.
The theoretical foundation of this approach stems from the observation that different quantization models exhibit distinct biases (unfairness) across various regions of the input data space Liu et al. (2025). Specifically, for certain subsets of data, one quantization model may consistently yield lower prediction errors (i.e., lower loss values) than others. For example, as shown in the Figure  2(a), we evaluate the perplexity of five mainstream quantization methods—GPTQ, SmoothQuant, K-Quants, imatrix, and AWQ on seven different sub-datasets (Subdata1–7) of the Qwen 1.7B model. The results show that there are significant differences in performance among these subsets. For instance, SmoothQuant has the lowest perplexity on Subdata4, while AWQ performs the best on Subdata2.
@@
By leveraging the router model to dynamically assign samples to the most suitable expert, MoQE effectively selects, at each inference step, the model with the lowest loss function value on this subset of data. As illustrated in the Figure  2(b), this strategy enables the system to synthesize a composite decision boundary that outperforms any individual quantization model. Consequently, the router model can be interpreted as a multi-class classifier, and its effective design is critical to ensuring that the overall system performance exceeds that of any constituent model.
@@
Figure 2: (a) The specific performance of multiple quantization methods for Qwen 1.7B on different sub-datasets (Subdata1 - 7) (b) The explanations of MoQE system which consists of two quantization experts. MoQE always selects the model with the lower loss on this subset of data.
@@
It is important to note that the router model operates on vectorized representations of input data, implying that raw data formats may first be transformed into appropriate embedding. This process is optional depending on the task. For example, in CV tasks, images are naturally represented as tensors and can be directly processed by the router model. Thus, CV tasks do not require embedding process. In contrast, for NLP tasks, textual inputs must be converted into dense vector representations—typically through embedding techniques—before they can be handled. The quality of this embedding process directly influences the router model’s classification accuracy.
@@
In the context of NLP, MoQE uses the pre-existing embedding layer of the original full-precision model as its embedding layer. This choice serves two purposes: (1) it avoids the need to train a separate embedding layer for routing, which would be computationally expensive and potentially degrade performance due to misalignment; and (2) it reduces inference cost because using the same embedding layer avoids an extra embedding pass. Therefore, the quantization experts’ embedding layer is left in full precision.
@@
In our experimental design, we primarily focus on quantization variants of a single full-precision model, reflecting typical industrial deployment scenarios where only one full-precision model is available for quantization. Moreover, in NLP tasks, the shared embedding layer constraint limits the framework to experts originating from the same base model, which implies that all quantization experts come from the same full-precision base model. However, the MoQE framework is, in principle, extensible to quantization models derived from multiple distinct full-precision models. Prior work Liu et al. (2025) suggests that for well-trained models, performance variations across quantization versions in different data regions are largely determined by local gradient and Hessian properties.
3.2 Router Model
@@
The design of router models must be tailored to the intrinsic characteristics of different task domains. For instance, NLP tasks necessitate a strong emphasis on contextual dependencies, while CV tasks primarily rely on spatial and hierarchical visual features. By carefully calibrating the architectural and parametric configurations of the router model, classification accuracy can be significantly improved, ensuring that each quantization expert model operates at peak efficacy within its targeted data distribution. This, in turn, contributes to enhanced overall inference accuracy. To this end, and to address the prevalent use cases in modern deep learning, we have designed two specialized router models tailored specifically for CV and NLP tasks, respectively—each optimized to exploit the structural priors inherent in its respective domain. This section details the router model’s design and implementation. Convergence and error analyses are provided in appendix A.5.
@@
We propose a lightweight, multi-head SEResNet-8 router architecture. The processing pipeline commences with a three-layer MLP tasked with the non-linear transformation of the input global features. Each linear transformation is followed by Batch Normalization (BN) and a ReLU activation. Subsequently, a three-stage SEResNet-8 module, configured with channel dimensions of 16, 32, and 64 and a stride pattern of 1→2→2, is employed to extract hierarchical local visual representations. Finally, to generate the routing logits, the resulting feature map is flattened into a sequence and fed into an 8-head self-attention module. This design is intended to aggregate global information, thereby facilitating an expert assignment strategy that is both globally-aware and load-balanced, all while ensuring minimal computational overhead.
NLP Router Model The core design principle of our NLP router is to endow router model decisions with the same level of semantic richness and contextual sensitivity as the expert computations.The structure of router model is shown in Figure  3.
Figure 3: Structure of the router model
The foundation of our architecture is a Transformer encoder. Its primary task is to process the entire input sequence in parallel, transforming each static token embedding into a context-aware representation. This initial stage captures the global linguistic environment surrounding each token. Building upon these general-purpose contextual vectors, a dedicated self-attention module subsequently performs a function of focused refinement. It distills and re-weights the existing representations to identify the most critical contextual cues, thereby optimizing the routing decision. This role forms a crucial complement to the general-purpose semantic understanding provided by the initial encoder. Finally, MLP maps the refined, context-rich feature vectors to the final routing logits. It employs a non-linear projection to translate these complex representations into definitive scores for each expert.
3.3 Router Fine-Tuning
In the fine-tuning stage of the MoQE system, the core objective is to train an efficient and stable router model that can dynamically assign input data to the most suitable quantization expert based on its features. To achieve this, a key principle is that the weights of all quantization experts remain completely frozen throughout the training process. The router model itself is maintained at full-precision during training to ensure it possesses maximum representational capacity for making accurate routing decisions.
Each sample is fed in parallel to all N frozen quantization experts, and the prediction loss is calculated for each expert. The expert that yields the lowest loss is designated as the ”optimal expert” for that sample, serving as the target label for the router model to learn. We designed a composite loss function to ensure that the router model not only learns to assign samples accurately but also avoids collapsing into a state where it relies on only a few experts. The total objective function consists of two components:
$\mathcal{L}=\mathcal{L}_{\text{CE}}+\alpha_{\text{dyn}}\mathcal{L}_{\text{bal}},\mathcal{L}_{\text{bal}}=N\cdot\sum_{i=1}^{N}P_{i}\cdot F_{i}$ (1)
$P_{i}\triangleq\mathbb{E}_{x\sim\mathcal{B}}\!\left[p_{i}(x)\right]\;\approx\;\frac{1}{B}\sum_{x\in\mathcal{B}}p_{i}(x),F_{i}\;\triangleq\;\frac{n_{i}}{B}\,$ (2)
Here, $P_{i}$ is the mean routing probability for expert $i$ and $F_{i}$: the fraction of samples actually dispatched to expert $i$. NN denotes the number of experts, $p_{i}(x)$ is the softmax probability of routing sample $x$ to expert $i$, B is the current mini-batch, $n_{i}$ is the number of samples in the batch actually dispatched to expert $i$. The influence of this balancing regularizer is controlled by a dynamic weight, $\alpha_{\text{dyn}}=\alpha(1+\sigma_{t})$, $t$ is the epoch number. Here, $\alpha$ is a hyper-parameter with an initial value of 0.02 , while $\sigma_{t}$ is the relative standard deviation of expert usage frequencies. We linearly decay $\alpha$ to 0 in the final stage of training, thereby shifting the optimization focus toward minimizing $\mathcal{L}_{\mathrm{CE}}$ and stabilizing the routing policy while preserving system diversity. Detailed training configurations are described in Appendix A.1.
4 Experiments